### v1

# GeoBenchX: Benchmarking LLMs for Multistep Geospatial Tasks

Abstract
In this paper, we establish a benchmark for evaluating large language models (LLMs) on multi-step geospatial tasks relevant to commercial GIS practitioners. We assess seven leading commercial LLMs (Sonnet 3.5 and 3.7, Haiku 3.5, Gemini 2.0, GPT-4o, GPT-4o mini, and o3-mini) using a simple tool-calling agent equipped with 23 geospatial functions. Our benchmark comprises tasks across four categories of increasing complexity, with both solvable and intentionally unsolvable tasks to test hallucination rejection. We develop an LLM-as-Judge evaluation framework to compare agent solutions against reference implementations. Results show Sonnet 3.5 and GPT-4o achieve the best overall performance, with Claude models excelling on solvable tasks while OpenAI models better identify unsolvable scenarios. We observe significant differences in token usage, with Anthropic models consuming substantially more tokens than competitors. Common errors include misunderstanding geometrical relationships, relying on outdated knowledge, and inefficient data manipulation. The resulting benchmark set, evaluation framework, and data generation pipeline are released as open-source resources, providing one more standardized method for ongoing evaluation of LLMs for GeoAI.
GeoBenchX: Benchmarking LLMs for Multistep Geospatial Tasks
Varvara Krechetova World Bank New York, USA                        Denis Kochedykov J.P.Morgan Chase New York, USA
1 Introduction
In this paper, we create a benchmark set and describe a simple LLM-as-Judge based methodology to evaluate the performance of several leading commercial large language models (LLMs) with function-calling abilities for solving a range of multi-step geospatial tasks. Geospatial tasks are distinct in that they require LLMs to have good factual geographical knowledge, understanding how to work with spatially defined data, georeferenced data, what a coordinate reference system is, understanding what are various types of spatial or geodata. We used a simple tool-calling agent to specifically assess the LLMs capabilities and have tested several models: Sonnet 3.5 and 3.7, Haiku 3.5, Gemini 2.0, GPT-4o, GPT-4o mini, o3-mini.
Our focus is on the commercial GIS practitioner requirements, which influence the choice of tasks, models and evaluation framework. Unlike other adjacent benchmarks, we focus on tasks that are closer to what GIS practitioners usually need for socio-economic analysis and mapping for reports, news articles, memos, and the like. Practitioners would use LLM equipped with GIS tools in a chat environment to get series of maps, numbers, and lists of locations. The tasks in the benchmark set reflect this – asking to make various types of map, to find geographical objects by some conditions, count objects within certain geographic boundaries, etc. Also, importantly for practitioners, we test the ability of LLMs to correctly recognize when the task cannot be solved with given tools and inform the user. This reduces the risk of hallucinated outputs. Next, we benchmark most current advanced commercial LLMs used in industry setting. One requirement of such industry usage is that the variability of outputs is not desired, so the models were tested at temperature 0. Lastly, the evaluation method we propose allows one to automate the ongoing evaluation which is increasingly important in practice with frequent models updates in commercial settings.
The benchmark set and code are publicly released at https://github.com/Solirinai/GeoBenchX.
The paper is structured as follows. First, we describe the task-solving agent and the set of tools/datasets. Second, we describe how the benchmark set of tasks was created and its composition. Third, we detail the approach to evaluation of the results of the task-solver agent using an LLM-as-Judge evaluator agent and testing the evaluator performance. Lastly, we compare the performance metrics of the solver agent based on different LLMs and the cost efficiency in terms of token usage.
2 Related works
Research in the field of GeoAI, a part of spatial data science at the intersection with Artificial Intelligence (AI) Janowicz et al. (2020), got a boost with recent advances in generative AI and the deployment of accessible LLMs. GeoAI researchers actively explore the capabilities of the available models in relation to geospatial data science, design their applications, and methods to improve their performance. The avenues explored include analysis of the ability of LLM (GPT-4) for spatial reasoning in travel networks, distance and elevation estimations, route and supply chain analysis Roberts et al. (2023); creating autonomous LLM-based GIS applications (LLM-Geo) Li and and (2023), GIS Copilot for GIS software Akinboyewa et al. (2024), ability LLMs to correctly understand and apply GIS toolsZhang et al. (2024), integrating LLM and cartography tools into MapGPT He et al. (2024), combining several LLMs to increase the efficiency of spatial data processing in the multi-agent framework for automated processing of vector data in shapefile format Lin et al. (2024), benefits of generative AI agents in smart mobilityXu et al. (2024), fine-tuning LLM to process natural language queries into executable codeMansourian and Oucheikh (2024). There is a benchmarking of LLM performance in spatial tasks in progress, with analysis of LLMs’ geospatial knowledgeBhandari et al. (2023); Renshaw et al. (2025), of GIS coding abilitiesGramacki et al. (2024), the recent study testing models on several groups of tasks that require geographic, GIS, and spatial knowledge, including changes in their performance in response to changes in prompts Liuchang Xu et al. (2025).
3 Agent, tools, and datasets
The experiment setup is based on the pre-built Langgraph ReAct agent LangChain AI (2024) equipped to call one of the selected LLMs with a list of attached tools. We intentionally choose a simple agent architecture to measure the impact of the LLMs on the final performance, not compensated and leveled by complex agent architectures.
The agent processed a given user task without Human-In-The-Loop involvement. Each task normally requires series of tool calls to solve and the agent does not always find the shortest solutions, so we allowed up to 25 iterations “LLM suggests tool call(s) – LLM receives the tool(s) response”. Less iterations were not enough for some complex tasks, while allowing more steps did not increase success rates. Experiments indicated that an increase in temperature leads to the agent being stuck looping over the same steps and to lower task completion rates, so, all LLMs with enabled temperature parameter were tested at temp. 0.
The agent system prompt sets the persona, briefly describes the setting, and the rules for solving the tasks (complete prompt in Appendix). The agent was instructed not to ask the user for clarifications and not to provide code to solve the task, but to respond with the resulting map, the names of the data sets used and explanations. The set of rules instructs when to reject the task and when to leave the solution empty (meaning that no tools are needed to answer user’s question), when to consider the data available (if the available data do not fully cover the geography asked in the task), and how to select columns for mapping data if the user did not specify (implied conditions frequent in production user questions).
The text of the task is supplied as the next message for the LLM Chat API as the user prompt.
Agent is provided with 23 tools (specifications are in the Appendix) that can read tabular data, geospatial data in vector or raster formats, perform operations on dataframes and geodataframes (merge, filter by categorical or numerical values), their columns, perform spatial operations on geodataframes and rasters (spatial joins, filtering by locations, creating buffers, generate contour lines from raster), visualize geodataframes, generate choropleth maps and heatmaps. The tools leverage GIS-specific libraries like GeoPandas den Bossche et al. (2024), Rasterio Gillies et al. (2013–), GDAL Rouault et al. (2025), Shapely Gillies et al. (2025), and common data analysis libraries like Matplotlib, Plotly, Pandas, and Numpy.
To distinguish the situations in which the agent stops solving the task because it decides that the task could not be solved with the provided data sets / tools, the agent was equipped with an additional tool rejecttask that simply returned the message that the task is unsolvable.
When a tool result included a dataframe or geodataframe, information about its length, columns, data types, and non-empty cells was included in the tool response message. Where dataframes or geodataframes needed to be passed between tools, they were stored in the data-store dictionary accessible to all tools.
For clarity, hereafter we refer to the set of tasks used to evaluate the agent as “benchmark set” and the statistical and geo data used by the tools to solve the tasks as “data sets”.
To make the experimentation consistent, we opted to store the snapshots of datasets in the local experimentation environment rather than fetching data from online APIs on every run. This allowed us to avoid variability of benchmarking due to possible data API connection issues, data API changes, and data changes behind the scenes.
Snapshots of data sets were organized as dictionaries with short descriptive titles as keys and filenames as values, one dictionary per type of data. When the tool designed to load the respective data was called, LLM had to choose the tool argument, the data set to load, from the list of the keys of the appropriate dictionary.
The data included: Statistical data: panel data in csv format (18 datasets): World Bank data on economic and population indicators, tuberculosis cases at subnational levels. Geometry data: vector data in shapefile format (21 datasets): polygon geometries (national or subnational units, lakes), points geometries (wildfire locations, urban places, Amtrak stations, earthquake occupancies, seaports, mineral extraction facilities, power stations), line geometries (rivers, railways). Raster data: (11 rasters): accumulated snow cover, flood extents, population data.
4 Benchmark set
To represent typical use by GIS professionals of a GIS assistant, tasks were single-sentence questions in natural language that might be asked by someone who needs a map, a number, a visualization for drafting an analytical research note, creating an action plan in international development, cross-checking information, etc. Some tasks were naturally ambiguous or missing information.
We created a set of 200+ tasks, grouped by type and complexity. Based on Wald’s method, this sample size corresponds approximately to the 5% margin of error for estimating a binary match rate with the 90% confidence. Each group included tasks that could be solved with the tools/datasets provided (solvable tasks), tasks that looked similar to solvable but could not be solved with the tools or datasets provided (unsolvable tasks). The groups of tasks are as follows:
Merge-visualize
(36 tasks): require to join tabular numeric data (panel data for one or multiple indicators, e.g. countries-by-years) with a geographic geometry data and make a choropleth or bivariate map. Example: Map the relationship between GDP per capita and electric power consumption per capita globally”. This is the most straightforward group, the tasks mostly require selection of correct datasets, and two (three in case of bivariate map) tool calls. Only 6 of these tasks were not solvable with given tools/data. Due to simplicity of the tasks, their content and solutions have limited variation; therefore, a limited number of them needed to have enough coverage.
Process-merge-visualize
(56 tasks): require some processing of either tabular or geometry data (filtering or operations on columns) and then creating a choropleth or bivariate map. Example: Compare rural population percentages in Western Asian countries”. The tasks in this group are more complicated, require the use of knowledge of the world geography and the use of a wider variety of tools. Half of the tasks, 28, are not solvable with given tools/data.
Spatial operations
(53 tasks): require spatial joins, creating buffers around selected geometries, checking overlaps between rasters, performing calculations on rasters, finding values from raster using geometry data or filtering geometries based on raster values, calculating distances. They might require a numerical answer, and the creation of actual maps might be optional. This group is more challenging than the previous two, as it requires the agent to work with different types of data set in one task, navigate the full set of available tools, and maintain the overall logic of the solution through all its steps. Example: How many people live within 1 km from a railway in Bangladesh?”. In this group, 23 of the tasks were not solvable.
Heatmaps, contour lines
(54 tasks): required performing spatial operations (selection, overlaps, etc.), understanding the contents of rasters, and making a heatmap or generating contour lines as one of the steps. Example: Create a heatmap of earthquake occurrences in regions with high population growth”. This was the most challenging set, as the agent had to figure out if the heatmaps and contour lines tools were actually able to solve the task and to make correct decisions on a complex set of arguments. In this group, 41 tasks ended up being unsolvable with available data/tools, which was a test for LLMs ability to recognize an unsolvable task and generate a rejection rather than a hallucinated answer.
4.1 Tasks creation
For each group, seed tasks were sourced from industry professionals and publications. Then more tasks were generated by using the Sonnet 3.5 model based on the seed tasks, the description of the tools, and the lists of available datasets. Then tasks were generated that should look similar to solvable tasks but could not be solved using the provided data and tools.
All the generated tasks were then manually reviewed to keep only high-quality tasks in the set.
4.2 Tasks annotation
Each of the 200+ tasks was annotated with one or multiple ground-truth reference solutions. To create the annotation, we have started by generating the draft reference solutions by the agent using GPT-4o or Sonnet 3.5 models. We then manually reviewed and rewrote every solution.
One reference solution could be (1) Sequence of tool calls with arguments; similar to a Python code snippet (2) Empty sequence [] of tool calls meaning that the task can be solved from the LLM’s world knowledge (3) A single special tool call reject_task() meaning that the task should be rejected as it cannot be solved with given tools/data.
A nonempty reference solution looked as follows for the evaluator:
#overall reference solution comment
#toolcall1 comment
toolcall1(arg11=..., arg12=..., ...)
#toolcall2 comment
toolcall2(arg21=..., arg12=..., ...)
Approximately 75% of the solvable tasks have reference solutions consisting of 4-6 steps (tool-calls) and  25% consist of more steps (up to 14 steps).
We used comments in annotations as a way to record information important for the evaluator agent to make a judgment about whether the candidate solution and the reference solution are semantically equivalent. For example, telling which variations in the solution or step are irrelevant or important.
One task sometimes allowed multiple solutions because tasks in practice are often ambiguous and multiple analytical approaches are possible for them. For example, task Compare the freshwater withdrawal between African countries with and without significant railway networks” can be solved: (1) by rejecting because it is not specific enough (user need to clarify what’s “signigicant” railway networks; (2) by making separate maps of railway networks and water withdrawals for African countries; (3) by combining the data and mapping bivariate map to show the relation; (4)by implying some threshold and computing statistical metrics for 2 groups of countries.
5 Evaluation approach
5.1 Matching-based evaluator
The 2 widely used evaluation approaches are execution-based and matching-based. However, comparing directly the outputs of the agent on geospatial analytical tasks would be quite challenging because the outputs contain multiple images/maps, tables, numbers, a narrative and sometime all of it as in the section 4.2 example. Annotation of such ground truth is not practical in this case, so we opted for matching-based approach. We used LLM-as-Judge with multiple LLM judges to compare the solutions generated by the agent (“candidate solutions”) and reference solutions for the same task. To leverage the LLMs ability to analyze code, we represented both reference and candidate solutions as snippets of Python code as in 4.2. We asked LLM judges to compare semantic equivalence of the solution; this took care of situations where reference and candidate solutions had different order of steps, different arguments for tool-calls or even took a different path to the same analytical outcome.
We also compared the lengths of the reference and candidate solutions, since the agent often takes some extra steps to figure out a path to answering the user question, and such extra steps are often costly.
For evaluation of the solutions, we used a panel of judges: Sonnet 3.5, GPT-4o, and Gemini 2.0. The evaluator prompt included: persona setting, description of all tools, scoring taxonomy, instructions on how evaluate solutions, examples of evaluated solutions, corresponding scores, and reasoning behind each score. The evaluator was instructed to analyze the candidate solution for the task and produce (1) reasoning for matching any of the reference solutions for the task and the candidate solution and (2) the matching score (‘2’: match, ‘0’: no match, ‘1’: partial match)
5.2 Evaluating the evaluator
We evaluated and tuned the evaluator itself on a set of 44 tasks randomly sampled with stratification of all tasks, respecting task groupings. We manually scored candidate solutions for these tasks, providing a matching score and reasoning for it. We then ran the evaluator on the same tasks candidate solutions and compared human annotator match scores, and the evaluator produced match scores.
Among the three models used as judges, the Sonnet 3.5 model matching scores were closest to human annotator scores. scoring (0.91 of the scores corresponded to manual scores), followed by GPT-4o (0.86) and Gemini 2.0 (0.82).
In the below results section, we use all the 3 above LLMs as judges and average their match frequency estimates.
6 LLMs evaluation results
The following models were tested in the solver agent: gpt-4o-2024-08-06, gpt-4o-mini-2024-07-18, o3-mini-2025-01-31, gemini-2.0-flas%h-001, claude-3-5-sonnet-20241022, claude-3-5-haiku-20241022, claude-3-7-sonne%t-20250219. Each LLM was tested on the same set of tasks, and all models were tested at temperature 0 (except o3-mini, which does not have a temperature parameter).
The frequency of each of the match scores was calculated (1) separately for the solvable tasks (104 tasks) and the unsolvable tasks (98 tasks) (2) for each group of tasks according to 4.
To mark unsolvable tasks, the agent was provided with a rejecttask tool that it could call either upfront or after some steps if it reached the conclusion that the task cannot be solved with available datasets and tools.
To benchmark how efficient an LLM is at solving the task compared to the reference solution, we calculated the mean difference between the length (number of steps) of the reference solution (or average length of reference solutions if multiple) and the length of candidate solution. We express this difference relative to the length of the reference solution (so 0.5 means 50% more steps in the candidate solution than the reference).
The first observation is that models that are more successful in identifying unsolvable tasks tend to be less successful in dealing with solvable tasks and vice versa (Table 1, Table 2). Anthropic models perform better on solvable tasks, but prioritize giving any answer at all over rejecting tasks, Sonnet 3.5 is the most balanced from the three in this aspect. Gemini 2.0 performs similarly on both kinds of tasks. Although the three OpenAI models are better at rejecting tasks than solving them, among them, the GPT-4o is well balanced. Overall, on solvable tasks, Sonnet 3.5 turned out to be the best model, followed by Sonnet 3.7, GPT-4o, Gemini 2.0 and Haiku 3.5 (Table 2). We can also see that for unsolvable tasks, the candidate solutions are much longer, indicating LLMs trying to find a solution and eventually giving up. For solvable tasks, candidate solutions are about the same length as reference ones, with notable exceptions of Sonnet 3.7 that tends to produce longer solutions than needed and o3-mini that tends to reject tasks more frequently than other models. The detailed statistics for each group of tasks are in tables 3, 4, 5, 6.
Model No Partial Match $\Delta$ in
match match solution
length
Sonnet 3.7 0.84 0.02 0.14 6.47
Sonnet 3.5 0.61 0.02 0.37 2.51
Haiku 3.5 0.79 0.01 0.20 4.39
GPT-4o 0.38 0.03 0.59 2.45
GPT-4o mini 0.37 0.02 0.62 2.71
o3-mini 0.23 0.02 0.76 0.20
Gemini 2.0 0.54 - 0.46 1.83
Table 1: Success rate in identifying the unsolvable tasks (98 tasks from total of 202)
Model No Partial Match $\Delta$ in
match match solution
length
Sonnet 3.7 0.27 0.23 0.50 0.39
Sonnet 3.5 0.28 0.19 0.53 -0.05
Haiku 3.5 0.34 0.23 0.43 0.03
GPT-4o 0.41 0.15 0.44 -0.01
GPT-4o mini 0.70 0.13 0.16 -0.03
o3-mini 0.69 0.10 0.21 -0.59
Gemini 2.0 0.43 0.14 0.43 -0.03
Table 2: Success rate in solvable tasks (104 tasks from total of 202)
Model No Partial Match $\Delta$ in
match match solutions
length
Sonnet 3.7 0.23 0.11 0.66 0.91
Sonnet 3.5 0.17 0.15 0.72 0.29
Haiku 3.5 0.25 0.21 0.54 0.51
GPT-4o 0.24 0.13 0.63 0.31
GPT-4o mini 0.49 0.19 0.32 0.31
o3-mini 0.50 0.14 0.36 -0.39
Gemini 2.0 0.25 0.06 0.68 0.16
Table 3: Success rate in solving ‘Merge-visualize’ tasks (36 tasks from total of 202)
Model No Partial Match $\Delta$ in
match match solution
length
Sonnet 3.7 0.50 0.14 0.36 2.37
Sonnet 3.5 0.30 0.10 0.60 0.40
Haiku 3.5 0.59 0.09 0.33 2.11
GPT-4o 0.26 0.09 0.65 0.29
GPT-4o mini 0.54 0.07 0.39 0.60
o3-mini 0.48 0.05 0.46 -0.38
Gemini 2.0 0.35 0.12 0.53 0.17
Table 4: Success rate in solving ‘Process-merge-visualize’ tasks (56 tasks from total of 202)
Model No Partial Match $\Delta$ in
match match solution
length
Sonnet 3.7 0.55 0.14 0.31 3.42
Sonnet 3.5 0.48 0.12 0.40 1.73
Haiku 3.5 0.50 0.17 0.33 2.49
GPT-4o 0.50 0.08 0.42 1.95
GPT-4o mini 0.69 0.05 0.26 2.44
o3-mini 0.55 0.07 0.39 -0.3
Gemini 2.0 0.63 0.07 0.30 2.07
Table 5: Success rate in solving ‘Spatial operations’ tasks (53 tasks from total of 202)
Model No Partial Match $\Delta$ in
match match solution
length
Sonnet 3.7 0.81 0.13 0.06 6.05
Sonnet 3.5 0.76 0.10 0.13 2.15
Haiku 3.5 0.82 0.06 0.12 3.06
GPT-4o 0.56 0.08 0.36 2.00
GPT-4o mini 0.44 0.03 0.53 1.62
o3-mini 0.39 - 0.61 0.17
Gemini 2.0 0.64 0.04 0.32 0.91
Table 6: Success rate in solving ‘Heatmaps, contour lines’ tasks (54 tasks from total of 202)
Model Input Output
tokens (mln) tokens (k)
Sonnet 3.7 24 269
Sonnet 3.5 12 157
Haiku 3.5 15 203
GPT-4o 7.5 57
GPT-4o mini 8 55
o3-mini 3.5 561
Gemini 2.0 7 40
Table 7: Tokens used for generation of solutions for the whole set
6.1 Error Analysis
Looking at cases where candidate solutions did not match reference, the main causes/types of errors were:
(1) Poor understanding of geometrical relations: mistaking centroids for full polygons, attempting to get centroids of lines, and not understanding that centroids can be outside object boundaries, etc, etc. (2) Outdated world knowledge – good for established international groupings but fails with recent changes (e.g., BRICS expansion in 2024) or ambiguous groupings (Upper/Lower Nile countries). (3) Over-reliance on world knowledge - using internally known spelling of names rather than dataset values for filtering, despite having tools to check column values. (4) Inefficient data manipulation – merging without pre-filtering (e.g., merging with all USA counties instead of first filtering to specific state). (5) Skipping geographic filtering steps – keeping continent-level data when only country-level was needed. (6) Attempting to solve with partial data, trying to be helpful by using whatever geography was available instead of rejecting tasks with insufficient data coverage. (7) Modifying tasks to make them solvable rather than solving the original problem.
7 Conclusions
We have created a benchmark dataset for multistep geo-spatial task completion, introduced an evaluation framework, and evaluated the leading mainstream commercial LLMs on these tasks. The task profile and the evaluation approach reflect the practical requirements of GIS professionals and the use of LLM in a GIS-assistant mode.
Experiments indicate that commercial LLMs with function calling can successfully solve geospatial tasks even with casual, ambiguous instructions. Performance varies by task complexity, spatial thinking requirements, and model helpfulness bias.
Sonnet 3.5 and GPT-4o showed the best overall results and a balance between identifying unsolvable tasks and solving solvable ones. Anthropic models performed well on easier tasks, but their preference for providing solutions rather than rejecting tasks, high token usage, and price may disadvantage them for geospatial tasks. o3-mini and GPT-4o-mini overall had lower success rates, especially in easy tasks, though they excel at identifying unsolvable tasks and are affordable.
Limitations
We did not tune prompts individually for each LLM; we evaluate only solution algorithm not the final response (for rejects and final numbers); the benchmark set size is moderate; we didn’t evaluate the variance of the responses (only the binomial confidence bounds).
Appendix A Prompts for solver agent
The system prompt is in Figure 1, the rules prompt is in Figure 2.
System Prompt You are a geographer who answers the questions with maps whenever possible. You do have a list of python functions that help you to load statistical data, geospatial data, rasters, merge dataframes, filter dataframes, get uniques values from columns, to plot a map, make spatial operations on raster and vector data. You do not provide code to generate a map, you provide names of the datasets you used; the statistical dataset, the geospatial dataset and the resulting map with legend and needed explanations. Today is {current_date}.
Figure 1: System prompt for solver agent.
Rules Prompt While solving the tasks, please, follow the next rules: - If a task is not a geospatial task, if it does not require calling provided tools and use provided datasets to solve, proceed with reply, no explanations needed, no need to call reject_task’. - If the task is a geospatial task, however, this task can NOT be solved with either available datasets or tools, call the tool reject_task’. - If user did not specify the date of the data to be plotted on the map, you map the latest available period that has non empty data for more than 70% of the objects in the datasets (example: if a question is about countries of Africa, you map the latest year that has data for more than 70% of countries in Africa). - For bivariate maps select the data for the same year for both variables. For instance, if for variable 1 the latest data are for 2023 and for variable 2 the latest data are for 2014, select data for 2014 for both variable 1 and variable 2. If the common year for which both datasets have data is 2012, take the data for 2012. - If available dataset or datasets cover only part of the requested geography, consider the data unavailable. - Do not ask questions to the user, proceed with solving the task till a result is achieved. When the information is missing try to infer the reasonable values.
Figure 2: Rules prompt for solver agent.
Appendix B GeoSpatial Tools List
loaddata
Loads statistical data from the catalog into a Pandas DataFrame.
  • dataset: Name of the dataset to load
  • outputdataframename: Name for the new DataFrame
loadgeodata
Loads vector geospatial data from the catalog into a GeoPandas GeoDataFrame.
  • geodataset: Name of the geodataset to load
  • outputgeodataframename: Name for the new GeoDataFrame
getrasterpath
Constructs path to raster data from the catalog.
  • rasterdataset: Name of the raster dataset to access
getrasterdescription
Gets description of a raster dataset including metadata and statistics.
  • rasterpath: Path to the raster file
  • outputvariablename: Optional name for storing results
analyzerasteroverlap
Analyzes overlap between two rasters and calculates statistics.
  • raster1path: Path to first raster file
  • raster2path: Path to second raster file
  • outputvariablename: Name for storing results
  • resamplingmethod1/2: Resampling methods
  • plotresult: Whether to display visualization
getvaluesfromrasterwithgeometries
Masks a raster using vector geometries and calculates statistics.
  • rasterpath: Path to the raster file
  • geodataframename: Name of GeoDataFrame with geometries
  • outputvariablename: Name for storing results
  • plotresult: Whether to display visualization
mergedataframes
Merges statistical and geospatial dataframes using key columns.
  • dataframename: Name of DataFrame with statistical data
  • geodataframename: Name of GeoDataFrame with spatial data
  • statkey: Column name in statistical data for joining
  • geokey: Column name in spatial data for joining
  • outputdataframename: Name for the merged DataFrame
getuniquevalues
Gets unique values from a column in a DataFrame/GeoDataFrame.
  • dataframename: Name of DataFrame to analyze
  • column: Column name to get unique values from
filtercategorical
Filters DataFrame/GeoDataFrame by categorical values.
  • dataframename: Name of DataFrame to filter
  • filters: Dictionary of column names and values
  • outputdataframename: Name for filtered DataFrame
filternumerical
Filters DataFrame/GeoDataFrame using numerical conditions.
  • dataframename: Name of DataFrame to filter
  • conditions: Query string with numerical conditions
  • outputdataframename: Name for filtered DataFrame
calculatecolumnstatistics
Calculates summary statistics for a numerical column.
  • dataframename: Name of DataFrame to analyze
  • columnname: Column to analyze
  • outputvariablename: Name for storing results
  • includequantiles: Whether to include standard quantiles
  • additionalquantiles: Additional quantiles to calculate
createbuffer
Creates buffer zones around geometries in a GeoDataFrame.
  • geodataframename: Name of source GeoDataFrame
  • buffersize: Size of the buffer in meters
  • outputgeodataframename: Name for result GeoDataFrame
makechoroplethmap
Creates a choropleth map visualization from a GeoDataFrame.
  • dataframename: Name of GeoDataFrame with map data
  • mappingkey: Column to visualize
  • legendtext: Legend title
  • colormap: Color scheme for the map
filterpointsbyrastervalues
Samples raster values at point locations and filters by threshold.
  • rasterpath: Path to the raster file
  • pointsgeodataframename: Name of GeoDataFrame with points
  • valuecolumn: Column to store raster values
  • outputgeodataframename: Name for filtered GeoDataFrame
  • filtertype: Type of threshold filter
  • threshold1: First threshold value
  • threshold2: Optional second threshold value
  • plotresult: Whether to display visualization
selectfeaturesbyspatialrelationship
Selects features based on spatial relationships.
  • featuresgeodataframename: GeoDataFrame to select from
  • referencegeodataframename: GeoDataFrame to select by
  • spatialpredicates: List of spatial relationships
  • outputgeodataframename: Name for result GeoDataFrame
  • plotresult: Whether to display visualization
calculatelinelengths
Calculates lengths of line features in kilometers.
  • geodataframename: Name of GeoDataFrame with line features
  • outputvariablename: Name for storing results
calculatecolumns
Performs operations between columns of DataFrames.
  • dataframe1name: Name of first DataFrame
  • column1name: Column from first DataFrame
  • dataframe2name: Name of second DataFrame
  • column2name: Column from second DataFrame
  • operation: Mathematical operation to perform
  • outputcolumnname: Name for results column
  • outputdataframename: Name for result DataFrame
scalecolumnbyvalue
Performs operations between a column and a value.
  • dataframename: Name of DataFrame to modify
  • columnname: Column to scale
  • operation: Mathematical operation
  • value: Numerical value for scaling
  • outputcolumnname: Name for results column
  • outputdataframename: Name for result DataFrame
makeheatmap
Creates an interactive heatmap from point data.
  • geodataframename: Name of GeoDataFrame with points
  • valuecolumn: Column for intensity values
  • outputhtmlname: Optional name for HTML output
  • centerlat/lon: Map center coordinates
  • zoomlevel: Initial zoom level
  • radius: Radius of influence for each point
  • mapstyle: Base map style
  • width/height: Map dimensions
visualizegeographies
Displays multiple GeoDataFrames as map layers.
  • geodataframenames: List of GeoDataFrame names
  • layerstyles: Optional styling for each layer
  • basemapstyle: Style of the basemap
  • title: Map title
  • addlegend: Whether to add a legend
getcentroids
Calculates centroids of polygon features.
  • geodataframename: Name of GeoDataFrame with polygons
  • outputgeodataframename: Name for centroid GeoDataFrame
plotcontourlines
Creates contour lines from raster data.
  • rasterpath: Path to the raster file
  • outputgeodataframename: Name for contour GeoDataFrame
  • interval: Interval between contour lines
  • min/maxvalue: Range for contours
  • plotresult: Whether to display plot
  • title: Plot title
generatecontoursdisplay
Creates contour lines using GDAL algorithm.
  • rasterpath: Path to the raster file
  • outputfilename: Name for output shapefile
  • contourinterval: Interval between contour lines
  • columntitle: Name for values column
  • nodataval: Value for no-data pixels
  • outputgeodataframename: Name for contour GeoDataFrame
  • min/maxvalue: Range for contours
  • plotresult: Whether to display visualization
makebivariatemap
Creates a bivariate choropleth map showing relationship between two variables.
  • dataframename: Name of GeoDataFrame with map data
  • var1/var2: Variable column names
  • var1name/var2name: Display names for variables
  • title: Map title
  • outputvariablename: Name for storing results
rejecttask
Returns a standardized message when task cannot be solved.